39 grams of an alloy of aluminium and magnesium when heated with excess of dilute hydrochloric acid form magnesium chloride , aluminium chloride and hydrogen. The evolved hydrogen collected at 0 degree centigrade has a volume of 44.8 litre at 1 atmospheric pressure calculate the composition of alloy by moles and there respected percentage

Grade:

2 Answers

Sunil Kumar FP
9 years ago
let the mole of magnesium be x and mole of aluminium be y
Mg + 2Hcl---MgCL2 +H2
Al +3Hcl ----AlCl3 +1.5H2
NOW
24x +27y=39
x +1.5y=2
solving these two equation we get
x=.5 ,y=1
percentage of magnesium=30.76%
percentage of lauminium=69.24%
Sunil Kumar FP
9 years ago
applying PV=nRT for hydrogen
n=2 mole
let the mole of magnesium be x and mole of aluminium be y
Mg + 2Hcl---MgCL2 +H2
Al +3Hcl ----AlCl3 +1.5H2
NOW
24x +27y=39
x +1.5y=2
solving these two equation we get x=.5 ,y=1 percentage of magnesium=30.76% percentage of lauminium=69.24%